A company needed ghc 1000 to finance its activities. The firm can financed this expenditure either by bonds or equity. Interest rate on bonds is 10%. The company can earn ghc 160 in good years and ghc80 in bad years. Assuming the firm faces equal probability of good and bad years;


(1) What will be the stream of returns on both bonds and equity if the company chooses the following financing options


a 100% equity financing

b 50% equity financing

c 20% equity financing

d 0% equity financing


(2) Estimate the equity risk associated with each option in (1)


(3) As an investor who wants to purchase a share in the company, which financing option will make you purchase the stock. Why????


1
Expert's answer
2021-09-04T15:18:12-0400

1) The overall return of the year will be:

"160*0.5+80*0.5=120"

the stream of returns on equity:

a) "120*100\%=120" per year

b) "120*50\%=60" per year

c) "120*20\%=24" per year

d) "120*0\%=0" per year

the stream of returns on bonds:

a) "120-1000*10\%*100\%=20" per year

b) "120-1000*10\%*50\%=70" per year

c) "120-1000*10\%*20\%=100" per year

d) "120-1000*10\%*0\%=120" per year

2) Self-financing provides more income, but if bad years last for a long period, the return on investments will be less than on bonds. However, if the project is financed with bonds, then the income will be small.

However, control over the company will be retained and equity capital can be invested in another financial instrument. But it will be more difficult to attract new funding later.

3) If the financing comes from equity capital, the shares will be sold. Since shareholder returns are dependent on the company's earnings, the bond option is not attractive because the stock returns will be too low because bond payments reduce earnings.